OpenStudy (anonymous):

Can someone please explain to me how to solve the linear system using elimination? I would appreciate it! Equation: 4x-y=5 4x+4y=-4

OpenStudy (anonymous):

you can multiply one of the equation by a negative so you can cancel out the 4x

OpenStudy (anonymous):

do you know how to do it?

OpenStudy (anonymous):

K ! In the system if we want ti find x or y first should chose one of them first and another in second . So should sum them and before that should do the work can delete one of the x s or y s . So should to make Symmetric from them . And in your equation find y is easyer caz we have 4x and 4x . So if we multiply one of the equations on -1 can have -4x and if sum -4x and 4x just have y . Look :(chose one of them to multiply on -1) -1(4x-y=5)=-4x+y=-5 4x+4y=-4 Sum them : -4x+4x+y+4y=-5-4 have : 5y=-9 y = -9/5 Got it ? @justintaylor By the way welcome in open study !

OpenStudy (anonymous):

Use the first equation to minus the second one in order to eliminate x. 4x-y-(4x+4y)=5-(-4) 4x-y-4x-4y=5+4 -5y=9 y=-9/5. Now we get y, so it is easy to get x. 4x-(-9/5)=5 4x=5-9/5 4x=25/5-9/5 4x=16/5 x=16/20 x=4/5 Thus, x=4/5 y=-9/5 If you have any more questions, I will be glad to answer them for you.

OpenStudy (anonymous):

take a look very detailed great examples 1) http://www.youtube.com/watch?v=y49duU5LeGQ 2) http://www.youtube.com/watch?v=6N73_lhZG9s

OpenStudy (campbell_st):

here is an easy method, look at the equations.... y is negative in the 1st and positive in the 2nd. so eliminate y multiply the 1st equation by 4 and add to the 2nd equation 16x -4y = 20 4x +4y = -4 -------------- 20x = 16 solve for x when you get x substitute into either equation to find y. always look at the equations and decide which is the easier letter to eliminate I chose y because of opposite signs while everyone else chose x... I'll let you decide hope this helps and makes sense